Therandomvariables X andY haveajointPDFgivenby {2e−2x fX,Y (x,y) = ComputeCov(X,Y). x , 0≤x<∞,0≤y≤x, 0, otherwise.

Therandom variables X andY have a joint PDF given by {fX,Y (x,y) =2e−2x /x   0≤x<∞,0≤y≤x, 0, otherwise. ComputeCov(X,Y). ,
Answer #1

-2x f(ay) = x ocazoo; olyca CoV (X4) = E(XV) - EQ EMY) fca) = 2 2 2x dy = 2 ē 2. E 22eux de Lot azz du= è 2x duz ax V = -le E

EXY) = 1) = ss xyze 2x dy doc s g2 e 227 a doc x2 e 2x do ret t=22 po tet at CON (x,y) = EXY) - GX ECY II) 24